# A theory of ecological invasions and its implications for eco-evolutionary dynamics

### v1

@@ Appendix D Derivation of linear response in CRMs
$X_{0}=\begin{pmatrix}\vec{N_{0}}\\ 0\end{pmatrix},\quad\vec{u_{0}}=\begin{pmatrix}-m_{0}\\ 0\end{pmatrix},\quad{A}_{0S/E}=\begin{pmatrix}0&-C_{0S/E}\end{pmatrix},\quad{A}_{S/E0}=\begin{pmatrix}0\\ E_{S/E0}^{\operatorname{T}}\end{pmatrix}$
@@
$N_{0}=\frac{g_{0}-C_{0\mathcal{E}}R_{\mathcal{E}}-{C}_{0\mathcal{S}}E_{\mathcal{S}\mathcal{S}}^{T}B^{-1}C_{\mathcal{S}\mathcal{E}}\vec{R}_{\mathcal{E}}+{C}_{0\mathcal{S}}PQ_{\mathcal{S}\mathcal{E}}\vec{N}_{\mathcal{E}}}{-C_{0\mathcal{S}}PE_{\mathcal{S}0}^{\operatorname{T}}}$ (54)
where invader fitness is $-m_{0}+C_{0\mathcal{E}}R_{\mathcal{E}}+C_{0\mathcal{S}}R_{\mathcal{S}}$ And the abundance shifts to consumers and resources are
$\begin{pmatrix}\delta N\\ \delta R\end{pmatrix}=\begin{pmatrix}B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}\\ P\end{pmatrix}E_{\mathcal{S}0}^{\operatorname{T}}N_{0}+\begin{pmatrix}B^{-1}&B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}\\ Q^{-1}_{\mathcal{S}\mathcal{S}}E^{\operatorname{T}}_{\mathcal{S}\mathcal{S}}B^{-1}&P\end{pmatrix}\begin{pmatrix}0&-C_{\mathcal{S}\mathcal{E}}\\ E^{\operatorname{T}}_{\mathcal{S}\mathcal{E}}&Q_{\mathcal{S}\mathcal{E}}\end{pmatrix}\begin{pmatrix}N_{\mathcal{E}}\\ R_{\mathcal{E}}\end{pmatrix}$ (55)
@@
$\delta N_{\mathcal{S}}$ $=B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}E_{\mathcal{S}0}^{\operatorname{T}}N_{0}-B^{-1}C_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}}+B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}(E^{\operatorname{T}}_{\mathcal{S}\mathcal{E}}N_{\mathcal{E}}+Q_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}})$ (56)
$\delta R_{\mathcal{S}}$ $=PE_{\mathcal{S}0}^{\operatorname{T}}N_{0}-Q^{-1}_{\mathcal{S}\mathcal{S}}E^{\operatorname{T}}_{\mathcal{S}\mathcal{S}}B^{-1}C_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}}+P(E^{\operatorname{T}}_{\mathcal{S}\mathcal{E}}N_{\mathcal{E}}+Q_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}})$ (57)
Appendix E Linear response of nonlinear consumer resource model
We can apply the linearization framework to specially non-linear consumer resource models, similarly to how we map general nonlinear ecological models to general Lotka-Volterra models.
Consider nonlinear consumer resource models
@@ Appendix E Linear response of nonlinear consumer resource model
00 0 $=g^{N}_{i}(\vec{N}+\delta\vec{N},\vec{R}+\delta\vec{R},\vec{m}+\delta\vec{m})$ (62)
00 0 $=g^{R}_{\alpha}(\vec{N}+\delta\vec{N},\vec{R}+\delta\vec{R},\vec{k}+\delta\vec{k})$ (63)
@@
Defining each susceptibility matrix,
@@
$\left.\frac{\partial g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})}{\partial N_{j}}\right|_{\vec{N}=\vec{N}_{\mathcal{S}\mathcal{S}}}=-\mathcal{E}_{\alpha j},\quad\left.\frac{\partial g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})}{\partial R_{\alpha}}\right|_{\vec{R}=\vec{R}_{\mathcal{S}\mathcal{S}}}=-\mathcal{Q}^{R}_{\alpha\beta},\quad\frac{\partial g^{R}_{\alpha}(\vec{X}_{\mathcal{S}\mathcal{S}},\vec{m}_{i})}{\partial{k_{\beta}}}\delta k_{\beta}=\mathcal{B}^{R}_{\alpha\beta}.$ (67)
@@
$\begin{pmatrix}Q^{N}&C\\ -E&Q^{R}\end{pmatrix}\begin{pmatrix}\delta\vec{N}\\ \delta\vec{R}\end{pmatrix}=\begin{pmatrix}\mathcal{B}^{N}&0\\ 0&\mathcal{B^{R}}\end{pmatrix}\begin{pmatrix}\delta\vec{m}\\ \delta\vec{k}\end{pmatrix}$ (70)
$\begin{pmatrix}(\mathcal{B}^{N})^{-1}&0\\ 0&(\mathcal{B^{R}})^{-1}\end{pmatrix}\begin{pmatrix}Q^{N}&C\\ -E&Q^{R}\end{pmatrix}\begin{pmatrix}\delta\vec{N}\\ \delta\vec{R}\end{pmatrix}=\begin{pmatrix}\delta\vec{m}\\ \delta\vec{k}\end{pmatrix}$ (71)
@@
Appendix F Linear response in the microbial consumer resource model
@@
00 0 $=\sum_{\alpha}(1-l_{\alpha})w_{\alpha}c_{i\alpha}R_{\alpha}-m_{i}$ (79)
00 0 $=\frac{K_{\alpha}}{\omega}-R_{\alpha}+\sum_{j}\frac{\kappa_{i\alpha}^{\text{eff}}}{\omega}N_{j}$ (80)
Applying the formulation in the last section to linearize nonlinear consumer resource model, the self-consistency equations for microbial consumer resource model can be obtained by mapping it to the linear consumer resource model in the last section
@@
Appendix G Algorithm to solve the self-consistency formula
Here we present a simple algorithm to solve the self-consistency formula iteratively. The code can be found at https://github.com/Emergent-Behaviors-in-Biology/invasion-theory
@@ Appendix G Algorithm to solve the self-consistency formula
1:Input: Pre-invasion system interaction matrix $A$, species abundances $\vec{X}$, growth rates $\vec{r}$, invader interaction matrices $A_{II}$, $A_{IS}$, $A_{SI}$, and invader growth rates $u_{I}$, maximum number of iterations max_num_iters.
@@
7:     Update extinction boolean $Ebool$:
@@
Appendix H Numerical details for simulations
H.1 Lotka-Volterra Model
In Fig. 2. (b)-(c), and Fig. 4(c)-(e), we simulated an ensemble of 640 Lotka-Volterra models. We assembled each system’s pre-invasion community with $n=10$ species, which resulted in an average of $8$ surviving species. The interaction matrices followed statistics
@@ H.1 Lotka-Volterra Model
And the environmental parameters similarly were independent and had mean $\mu_{r}/n$ and variance $\sigma_{r}^{2}/n$. In this figure we chose $\mu_{A}=2$, $\sigma_{A}=0.2$, $\rho=1$ for a symmetric model, and $\mu_{r}=2$, $\sigma_{r}=0.1$. We chose the threshold for extinction as $10^{-5}$. Invaders were uncorrelated to the system but shared the same statistics of interaction matrix and environmental parameters. We performed simulations with numerical ODE simulator ’LSODA’, with hyperparameters atol=1e-12, rtol=1e-12, and checked convergence by ensuring the maximum derivative was less than 1e-12.
H.2 MacArthur Consumer Resource model
In Fig. 2(e)-(g), we simulated an ensemble of 640 generalized MacArthur consumer resource models. We assembled each system’s pre-invasion community with $60$ consumers and $n=30$ resources, which resulted in an average of 7 surviving consumers and 22 surviving resources. In this model, we set consumer extinction threshold at $10^{-3}$ and resource extinction threshold at $10^{-9}$. The consumer preference matrix followed statistics:
$\langle C_{i\alpha}\rangle=\frac{\mu_{C}}{n},\langle C_{i\alpha}C_{j\beta}\rangle=\delta_{i\alpha}\delta_{j\beta}\frac{\sigma_{C}^{2}}{n}$
@@ H.2 MacArthur Consumer Resource model
$\langle Q_{ij}\rangle=\mu_{Q}/n+\delta_{ij}$
$\langle Q_{ij}Q_{kl}\rangle=\delta_{ij}\delta_{kl}\frac{\sigma_{Q}^{2}}{n}+\delta_{il}\delta_{jk}\frac{\rho_{Q}\sigma_{Q}^{2}}{n},$
@@
For environmental parameters, we had mean and variance of consumer mortality rate, $\mu_{m}/n$, $\sigma_{m}^{2}/n$, with $\mu_{m}=1$, $\sigma_{m}=0.1$. And mean and variance of carrying capacity were, $\mu_{k}/n$, $\sigma_{k}^{2}/n$, with $\mu_{k}=2$, $\sigma_{k}=0.1$. Again, the invader parameters were uncorrelated with the system and followed the same statistics. We used a constraint optimization algorithm instead of numerical ODE for simulation. The algorithm was similar to the one used in community simulator [cite].
H.3 Microbial consumer resource model
In Fig. 2 (i)-(k), Fig. 4 (b)-(e) with legend “Assembled”, and Fig. 3 (f)-(i), we simulated an ensemble of 640 microbial consumer resource models (MiCRM). We assembled each system’s pre-invasion community with $60$ consumers and $n=30$ resources, which resulted in an average of 8 surviving consumers, while resources were abiotic and did not go extinct. It had a cross-feeding structure where we supplied one of the resources at a high rate $k=1000$, while others were by-products. Similar to before, the consumer preference matrix was
$\langle c_{i\alpha}\rangle=\frac{\mu_{c}}{n},\langle c_{i\alpha}c_{j\beta}\rangle=\delta_{i\alpha}\delta_{j\beta}\frac{\sigma_{c}^{2}}{n}$
@@ H.3 Microbial consumer resource model
H.4 Eco-evolutionary dynamics of MiCRM

### v2

@@ Appendix E Derivation of linear response in CRMs
$X_{0}=\begin{pmatrix}\vec{N_{0}}\\ 0\end{pmatrix},\quad\vec{r_{0}}=\begin{pmatrix}-m_{0}\\ 0\end{pmatrix},\quad{A}_{0S/E}=\begin{pmatrix}0&-C_{0S/E}\end{pmatrix},\quad{A}_{S/E0}=\begin{pmatrix}0\\ E_{S/E0}^{\operatorname{T}}\end{pmatrix}$ (58)
@@
$N_{0}=\frac{f_{0}-C_{0\mathcal{E}}R_{\mathcal{E}}-{C}_{0\mathcal{S}}E_{\mathcal{S}\mathcal{S}}^{T}B^{-1}C_{\mathcal{S}\mathcal{E}}\vec{R}_{\mathcal{E}}+{C}_{0\mathcal{S}}PQ_{\mathcal{S}\mathcal{E}}\vec{N}_{\mathcal{E}}}{C_{0\mathcal{S}}(1-P)Q^{-1}_{\mathcal{S}\mathcal{S}}E_{\mathcal{S}0}^{\operatorname{T}}}$ (62)
where invader fitness is $f_{0}=C_{0\mathcal{E}}R_{\mathcal{E}}+C_{0\mathcal{S}}R_{\mathcal{S}}-m_{0}$ And the abundance shifts to consumers and resources are
$\begin{pmatrix}\delta N_{\mathcal{S}}\\ \delta R_{\mathcal{S}}\end{pmatrix}$ $=\begin{pmatrix}B^{-1}&B^{-1}CQ^{-1}\\ Q^{-1}E^{\operatorname{T}}B^{-1}&(I-P)Q^{-1}\end{pmatrix}_{\mathcal{S}\mathcal{S}}\left[\begin{pmatrix}0&-C\\ E^{\operatorname{T}}&Q\end{pmatrix}_{\mathcal{S}\mathcal{E}}\begin{pmatrix}N_{\mathcal{E}}\\ R_{\mathcal{E}}\end{pmatrix}-\begin{pmatrix}0&-C\\ E^{\operatorname{T}}&Q\end{pmatrix}_{\mathcal{S}0}\begin{pmatrix}N_{0}\\ 0\end{pmatrix}\right$] (63)
$=\begin{pmatrix}B^{-1}&B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}\\ Q^{-1}_{\mathcal{S}\mathcal{S}}E^{\operatorname{T}}_{\mathcal{S}\mathcal{S}}B^{-1}&(I-P)Q^{-1}_{\mathcal{S}\mathcal{S}}\end{pmatrix}\begin{pmatrix}-C_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}}\\ E^{T}_{\mathcal{S}\mathcal{E}}N_{\mathcal{E}}+Q_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}}-E^{T}_{\mathcal{S}0}N_{0}\end{pmatrix}$ (64)
@@
$\delta N_{\mathcal{S}}$ $=-B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}E_{\mathcal{S}0}^{\operatorname{T}}N_{0}-B^{-1}C_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}}+B^{-1}C_{\mathcal{S}\mathcal{S}}Q_{\mathcal{S}\mathcal{S}}^{-1}(E^{\operatorname{T}}_{\mathcal{S}\mathcal{E}}N_{\mathcal{E}}+Q_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}})$ (65)
$\delta R_{\mathcal{S}}$ $=-(I-P)Q^{-1}_{\mathcal{S}\mathcal{S}}E_{\mathcal{S}0}^{\operatorname{T}}N_{0}-Q^{-1}_{\mathcal{S}\mathcal{S}}E^{\operatorname{T}}_{\mathcal{S}\mathcal{S}}B^{-1}C_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}}+(I-P)Q^{-1}_{\mathcal{S}\mathcal{S}}(E^{\operatorname{T}}_{\mathcal{S}\mathcal{E}}N_{\mathcal{E}}+Q_{\mathcal{S}\mathcal{E}}R_{\mathcal{E}})$ (66)
Appendix F Linear response of nonlinear consumer resource model
We can apply the linearization framework to specially non-linear consumer resource models, similarly to how we map general nonlinear ecological models to generalized Lotka-Volterra models.
Consider general consumer resource models of the form
@@ Appendix F Linear response of nonlinear consumer resource model
$0$ $=g^{N}_{i}(\vec{N}+\delta\vec{N},\vec{R}+\delta\vec{R},\vec{m}+\delta\vec{m})$ (71)
$0$ $=g^{R}_{\alpha}(\vec{N}+\delta\vec{N},\vec{R}+\delta\vec{R},\vec{k}+\delta\vec{k})$ (72)
@@
Defining each effective interaction matrix,
@@
$\left.\frac{\partial g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})}{\partial N_{j}}\right|_{\vec{N}=\vec{N}_{\mathcal{S}\mathcal{S}}}=-\mathcal{E}_{\alpha j},\quad\left.\frac{\partial g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})}{\partial R_{\beta}}\right|_{\vec{R}=\vec{R}_{\mathcal{S}\mathcal{S}}}=-\mathcal{Q}^{R}_{\alpha\beta},\quad\frac{\partial g^{R}_{\alpha}(\vec{X}_{\mathcal{S}\mathcal{S}},\vec{m}_{i})}{\partial{k_{\beta}}}\delta k_{\beta}=\mathcal{B}^{R}_{\alpha\beta}.$ (76)
@@
$\begin{pmatrix}Q^{N}&\mathcal{C}\\ -\mathcal{E}&Q^{R}\end{pmatrix}\begin{pmatrix}\delta\vec{N}\\ \delta\vec{R}\end{pmatrix}=\begin{pmatrix}\mathcal{B}^{N}&0\\ 0&\mathcal{B^{R}}\end{pmatrix}\begin{pmatrix}\delta\vec{m}\\ \delta\vec{k}\end{pmatrix}$ (79)
$\begin{pmatrix}(\mathcal{B}^{N})^{-1}&0\\ 0&(\mathcal{B^{R}})^{-1}\end{pmatrix}\begin{pmatrix}Q^{N}&\mathcal{C}\\ -\mathcal{E}&Q^{R}\end{pmatrix}\begin{pmatrix}\delta\vec{N}\\ \delta\vec{R}\end{pmatrix}=\begin{pmatrix}\delta\vec{m}\\ \delta\vec{k}\end{pmatrix}$ (80)
Figure S3: Dressed invasion fitness and extinction-induced feedbacks govern invasion outcomes in additional consumer-resource models. Simulations of (a–d) the MacArthur Consumer Resource Model (MCRM) and (e–h) a Monod growth consumer-resource model show qualitatively similar results as the Microbial Consumer Resource Model (MiCRM) presented in the main text (Fig.3).
@@
Appendix G Linear reponse in Monod CRM with type II functional response
This section explicitly demonstrates the applicability of our framework to ecological models of consumer-resource dynamics where the effective interactions depend on community state as described in last section. One generic way this can occur is when the per capita growth rate functions $g_{i}^{N}(\vec{N},\vec{R},\vec{m})$ of consumer species $i$ are nonlinear functions of consumer abundances $N_{i}$ and/or resource abundances $R_{\alpha}$ (see Eqs. (67) and (68) for definition).
A paradigmatic case of such consumer-resource models is an extension of the MacArthur consumer-resource model (where $g_{i}$ is a linear function, even though the full dynamics are nonlinear). In this extension, consumer species dynamics $\dot{N_{i}}$ are governed by Monod growth rate functions (Type II functional responses) instead of linear functions of $R_{\alpha}$, as
$\frac{dN_{i}}{dt}$ $=N_{i}\left(\sum_{\alpha}\frac{C_{i\alpha}R_{\alpha}}{h+R_{\alpha}}-m_{i}\right),$ (85)
$\frac{dR_{\alpha}}{dt}$ $=R_{\alpha}\left(K_{\alpha}-R_{\alpha}-\sum_{j}\frac{E_{j\alpha}N_{j}}{h+R_{\alpha}}\right).$ (86)
The steady state condition is
$0$ $=g^{N}_{i}(\vec{N},\vec{R},\vec{m})=\sum_{\alpha}\frac{C_{i\alpha}R_{\alpha}}{h+R_{\alpha}}-m_{i},$ (87)
$0$ $=g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})=K_{\alpha}-R_{\alpha}-\sum_{j}\frac{E_{j\alpha}N_{j}}{h+R_{\alpha}}.$ (88)
Effective interactions. The main difference between Type I and Type II responses is that the consumer preferences $C_{i\alpha}$, impacts $E_{i\alpha}$, and $Q_{\alpha\beta}$ must now be replaced by effective quantities $\mathcal{C}_{i\alpha}$, $\mathcal{E}_{\alpha j}$, and $Q_{\alpha\beta}^{(R)}$ that have to be calculated from the full non-linear dynamics. These effective quantities are precisely what we call effective interactions, since they represent how consumer species interact with resources and vice versa. Note also that $Q_{\alpha\beta}^{(R)}$ captures the effective interactions between resources. More generic models might similarly have effective interactions between species, say $B_{ij}^{(N)}$, but here $B=0$. One therefore has
$\mathcal{C}_{i\alpha}=\quad\left.\frac{\partial g^{N}_{i}(\vec{N},\vec{R},\vec{m})}{\partial R_{\alpha}}\right|_{\vec{R}=\vec{R}_{\mathcal{S}\mathcal{S}}}=\frac{C_{i\alpha}h}{(h+R_{\alpha})^{2}},$ (89)
$\mathcal{Q}^{R}_{\alpha\beta}=-\left.\frac{\partial g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})}{\partial R_{\beta}}\right|_{\vec{R}=\vec{R}_{\mathcal{S}\mathcal{S}}}=\delta_{\alpha\beta}-\delta_{\alpha\beta}\sum_{j}\frac{E_{j\alpha}N_{j}}{(h+R_{\alpha})^{2}},$ (90)
and
$\mathcal{E}_{\alpha j}=-\left.\frac{\partial g^{R}_{\alpha}(\vec{N},\vec{R},\vec{m})}{\partial N_{j}}\right|_{\vec{N}=\vec{N}_{\mathcal{S}\mathcal{S}}}=\frac{E_{j\alpha}}{h+R_{\alpha}}.$ (91)
Using these effective interactions, we can subsequently solve for the invader abundance in the same way as performed with the MCRM and MiCRM models. The remaining challenge lies in addressing the nonlinear invasion fitness:
$f_{0}=\sum_{\alpha}\frac{C_{0\alpha}R_{\alpha}^{\text{old}}}{h+R_{\alpha}}-m$ (92)
For convenience, in this model, this nonlinearity can be absorbed in the previous expression of invasion fitness simply by $C_{0\alpha}\to\frac{C_{0\alpha}}{h+R_{\alpha}}$. In summary, there are 4 replacements to normal MCRM solutions, including
$C_{i\alpha}$ $\to\frac{C_{i\alpha}h}{(h+R_{\alpha})^{2}},$ (93)
$E_{j\alpha}$ $\to\frac{E_{j\alpha}}{(h+R_{\alpha})},$ (94)
$Q_{\alpha\beta}$ $\to\delta_{\alpha\beta}-\delta_{\alpha\beta}\sum_{j}\frac{E_{j\alpha}N_{j}}{(h+R_{\alpha})^{2}},$ (95)
$C_{0\alpha}$ $\to\frac{C_{0\alpha}}{h+R_{\alpha}}\text{, \quad for invasion fitness only},$ (96)
Since all effective interaction matrices are proportional to $1/(h+R_{\alpha})$, changes in this quantity after invasion serve as a good measure of the model’s nonlinearity. In Fig. S5, we demonstrate that the mean relative error of the framework increases linearly with the mean relative shift of this value.
Appendix H Linear response in the microbial consumer resource model
@@
$0$ $=\sum_{\alpha}(1-l_{\alpha})w_{\alpha}c_{i\alpha}R_{\alpha}-m_{i}$ (100)
$0$ $=\frac{K_{\alpha}}{\omega}-R_{\alpha}+\sum_{j}\frac{\kappa_{i\alpha}^{\text{eff}}}{\omega}N_{j}$ (101)
Applying the formulation in the last section to arrive at the effective interactions in this Microbial consumer resource model, we arrive at the following effective interactions by following the same steps as in the last two appendices:
@@
Appendix I Algorithm to solve the self-consistency equations
Here we present a simple algorithm to solve the self-consistency equations iteratively. Algorithm 1 outlines the basic logic and implementation steps. We assess convergence by monitoring the extinction boolean indicators that determine the extinction and survival status of each species. The iteration terminates when these indicators remain unchanged following an update, indicating that the extinction and survival states have stabilized. If convergence is not reached, we impose a maximum of 50 iterations, after which we accept the last extinction boolean prediction as the final result. While Algorithm 1 performs reasonably well on its own, we incorporate several numerical improvements to enhance stability and better accommodate complexities arising in simulation results and experimental data:
First, when calculating the abundance shifts of resident species, we introduce a momentum hyperparameter at step 6. We chose its value to be $0.1$, meaning the new iteration contributes $10\%$ weight while the accumulated value carries $90\%$ weight. This parameter incorporates a weighted contribution from the abundance shift of the previous iteration into the current update, smoothing fluctuations and accelerating convergence. Such momentum-based updates are a well-known technique in numerical optimization to improve stability. However, to truly reflect the self-consistency equations, we use the final extinction boolean to recompute the abundances of both invaders and surviving species, so the memory from previous iterations is only retained in this boolean state.
Second, numerical uncertainties in abundance data before invasion mean that extinct species may not reach exact zero abundances in practice. To address this, at step 7 we implement a survival threshold, a small cutoff value ranging from $10^{-3}$ to $10^{-7}$ depending on the magnitude of species abundances in the particular model. Species with predicted abundances below this threshold are treated as extinct, instead of below exactly zero as described in step 7. This step effectively handles numerical noise and improves extinction boolean predictions.
Third, for models with nonlinear invasion fitness terms, such as Monod consumer resource model, we provide an additional input argument representing the effective consumer preference matrix used to compute nonlinear invasion fitness. This matrix is distinct from the ordinary interaction matrix $A$ and extend to complex models (see Appendix G for the detail example for Monod model).
Together, these modifications improve the robustness and applicability of the iterative algorithm across diverse systems, allowing accurate prediction of species abundances and extinction patterns following invasion.
The full code implementation underlying all results in this paper is publicly available at https://github.com/Emergent-Behaviors-in-Biology/invasion-theory
@@ Appendix I Algorithm to solve the self-consistency equations
1:Input: Pre-invasion system interaction matrix AA, species abundances $\vec{X}$, growth rates $\vec{r}$, invader interaction matrices $A_{II}$, $A_{IS}$, $A_{SI}$, and invader growth rates $u_{I}$, maximum number of iterations max_num_iters.
@@
7:  Update extinction boolean EboolEbool:
@@
Appendix J Error analysis
Figure S4: Error decreases with system size With the same simulation data used in Fig.2, we grouped predictions into different number of consumers after the invasions. For (a) Lokta-Volterra and (b) MacArthur consumer resource models, the decreasing trend saturate when the system is large enough because the self-consistency equation is exact and errors are only coming from the algorithm to deduce identity of extinct species, which also improves with system size. For (c) consumer resource models with Monod functional response and (d) Microbial consumer resource models, the error shows consistent decreasing trend because of the linear-response approximation to non-linear models.
The errors observed in the predictions compared to simulation results in Fig. 2 originate from multiple sources. Broadly, the error of the framework prediction arise from two components: (1) the self-consistency equations themselves and (2) the algorithm used to solve them. The primary source of error may vary depending on the particular model. Theoretically, the self-consistency equations are exact for models with linear steady-state conditions, including GLV Eq. (5), Eq. (6) and MCRM Eq.  (62) to (66) . To demonstrate this, Fig. S6 presents the results of solving the equations for the MCRM model using the same simulations as in Fig. 2(e)–(g), but with the extinction and survival states of each species after invasion provided explicitly rather than inferred through our iterative algorithm. This is done to disambiguate errors in our algorithm and approximations in our self-consistency equations. Notably, all visible errors in Fig. 2(e)–(g) disappear once the error introduced by the algorithm is removed (except numerical errors coming from ODE simulations themselves, between a relative $10^{-9}$ to $10^{-2}$. Since our iterative algorithm explores the boolean space of extinction and survival states locally, beginning from an initial condition where all resident species are assumed to survive, its performance improves when the true states do not deviate substantially from this starting point. Consequently, the mean relative error of invader abundance in Fig. S4(b) declines with number of consumers after the invasion, despite the error distribution remaining bimodal, with peaks centered around $10^{-7}$ (likely due to numerical simulation uncertainty) and $10^{-2}$ (attributable to the algorithm) respectively. A similar trend is observed in the Lotka-Volterra model shown in Fig. S4(a), although only two instances of substantial error occur, both at small consumer numbers of 4 and 5.
For nonlinear models such as MiCRM and Monod, the primary source of error likely arises from the self-consistency equations using a linear response approximation. This approximation assumes that the effective interactions among surviving species do not change significantly after invasion as described from Eq.(37) to (40). However, since these effective interactions depend on species or resource abundances, the resulting solutions from the self-consistency equations deviate from the actual invasion outcomes. This effect is clearest to test in the Monod model. As demonstrated in Fig. S5, the mean of the logarithm of the relative error increases nearly linearly with the mean of the logarithm of relative change in the effective interactions, confirming that deviations in these interactions contribute directly to the error. Moreover, this linear response approximation becomes more accurate in more species-diverse ecosystems, as the relative shift in species abundance caused by a single invader diminishes with increasing species number. This consistent trend is also evident in Fig. S4 (c)–(d), where the logarithm of the relative error decreases as the species diversity (number of surviving consumer species) increases.
Figure S5: Error increases with degree of nonlinearity in Monod consumer resource model The effective interaction in Monod consumer resource model are all proportional to $1/(h+R)$, The mean relative shift in this quantity serves as a direct measure of nonlinearity, quantifying the extent to which our linear-response approximation breaks down as nonlinearity increases.
Figure S6: Perfect prediction of MCRM when correct information of extinction is given. We demonstrate that the error distribution in Fig. 2(e)–(g) only has a larger second peak because of wrong boolean prediction of extinctions of species. When this information is given, the larger peak disappear and left with the smaller peak that only comes from numerical uncertainty of simulations.
Appendix K Numerical details for simulations
K.1 Lotka-Volterra Model
In Fig. 2(b)-(c), and Fig. 3(c)-(e), we simulated an ensemble of 640 Lotka-Volterra models. We assembled each system’s pre-invasion community with $n=10$ species, which resulted in an average of $8$ surviving species. The interaction matrices followed statistics
@@
And the environmental parameters similarly were independent and had mean $\mu_{r}/n$ and variance $\sigma_{r}^{2}/n$. In this figure we chose $\mu_{A}=2$, $\sigma_{A}=0.2$, $\rho=1$ for a symmetric model, and $\mu_{r}=2$, $\sigma_{r}=0.1$. We chose the threshold for extinction as $10^{-5}$. Invaders were uncorrelated to the system but shared the same statistics of interaction matrix and environmental parameters. We performed simulations with numerical ODE simulator ’LSODA’, with parameters atol=1e-12, rtol=1e-12, and checked convergence by ensuring the maximum derivative was less than 1e-12.
K.2 MacArthur consumer resource model
In Fig.  2(e)–(g) and Fig. S3(a)–(d), we simulated an ensemble of 640 generalized MacArthur consumer resource models. We assembled each system’s pre-invasion community with $60$ consumers and $n=30$ resources, which resulted in an average of 7 surviving consumers and 22 surviving resources. In this model, we set consumer extinction threshold at $10^{-3}$ and resource extinction threshold at $10^{-9}$. The consumer preference matrix followed statistics:
$\langle C_{i\alpha}\rangle=\frac{\mu_{C}}{n},\langle C_{i\alpha}C_{j\beta}\rangle=\delta_{i\alpha}\delta_{j\beta}\frac{\sigma_{C}^{2}}{n}$ (107)
@@
$\langle Q_{ij}\rangle=\mu_{Q}/n+\delta_{ij}$ (108)
$\langle Q_{ij}Q_{kl}\rangle=\delta_{ij}\delta_{kl}\frac{\sigma_{Q}^{2}}{n}+\delta_{il}\delta_{jk}\frac{\rho_{Q}\sigma_{Q}^{2}}{n},$ (109)
@@
For environmental parameters, we had mean and variance of consumer mortality rate, $\mu_{m}/n$, $\sigma_{m}^{2}/n$, with $\mu_{m}=1$, $\sigma_{m}=0.1$. And mean and variance of carrying capacity were, $\mu_{k}/n$, $\sigma_{k}^{2}/n$, with $\mu_{k}=2$, $\sigma_{k}=0.1$. Again, the invader parameters were uncorrelated with the system and followed the same statistics. We used a constraint optimization algorithm instead of numerical ODE for simulation. The algorithm was similar to the one used in [55].
K.3 Microbial consumer resource model
In Fig. 2 (m)-(o), Fig. 3 (f)-(i), and Fig. 4 (b)-(e) with legend “Assembled”, we simulated an ensemble of 640 microbial consumer resource models (MiCRM). We assembled each system’s pre-invasion community with $60$ consumers and $n=30$ resources, which resulted in an average of 8 surviving consumers, while resources were abiotic and did not go extinct. It had a cross-feeding structure where we supplied one of the resources at a high rate $k=1000$, while others were by-products. Similar to before, the consumer preference matrix was
$\langle c_{i\alpha}\rangle=\frac{\mu_{c}}{n},\langle c_{i\alpha}c_{j\beta}\rangle=\delta_{i\alpha}\delta_{j\beta}\frac{\sigma_{c}^{2}}{n}$ (110)
@@
K.4 Eco-evolutionary dynamics of MiCRM